OpenStudy (anonymous):

Medal! Math help 5x+3= 42

OpenStudy (photon336):

You're first step subtract three from both sides 5x+3 = 42 -3 -3 divide both sides by five 5x = 39 5 5 x = (39/5)

OpenStudy (photon336):

@Buttercup214 does this make sense?

OpenStudy (anonymous):

thank you so much! can you help me with more to see if im doing it right

OpenStudy (photon336):

Yes, absolutely

OpenStudy (anonymous):

ok im going to open up another one so i can give you another medal